OpenStudy (milanitaly1618):

Raj paid $31.86 for a new basketball. The price he paid included sales tax of 6.25%. Which equation can be used to find the price, p, of the basketball before tax? A. p + 0.0625p = 31.86 B. 6.25p = 31.86 C. 0.0625p = 31.86 D. p + 6.25p = 31.86

OpenStudy (jdoe0001):

\(\Large {\color{brown}{ a}}\%\ of\ b \implies \cfrac{{\color{brown}{ a}}}{100}\cdot b\implies {\color{brown}{ a}}\div 100\cdot b\) what would THAT make the 6.25% of price "p" then?

OpenStudy (milanitaly1618):

I think 0.0625

rebeccaxhawaii (rebeccaxhawaii):

any ideas?

OpenStudy (jdoe0001):

0.0625 is correct :) since 6.25/100 = 0.0625 so.. now, look at your choices

OpenStudy (anonymous):

i think is a

OpenStudy (anonymous):

answer is a

OpenStudy (jdoe0001):

if the basketball cost $31.86, including tax that means, the price of the ball is really "p" plus tax but the tax is 6.25% of 0.0625p so, add those two fellows, and you'd get the actual cost of the ball

OpenStudy (milanitaly1618):

I agree A

OpenStudy (milanitaly1618):

Is that correct?

OpenStudy (milanitaly1618):

?

OpenStudy (milanitaly1618):

Nevermind